OpenStudy (anonymous):

Find the circular permutation for 5 things taken 5 at a time. 120 24 1 <---- I'm thinking this is the answer, but I wanna confirm this.

jimthompson5910 (jim_thompson5910):

This page might help http://mathworld.wolfram.com/CircularPermutation.html

jimthompson5910 (jim_thompson5910):

If not, then let me know

OpenStudy (anonymous):

Following the formula, does this mean the answer is 24? Because that's what I got. @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yep, it's 24 since (n-1)! = (5-1)! = 4! = 24

OpenStudy (anonymous):

Thought so! Thank you.

jimthompson5910 (jim_thompson5910):

yw